OpenStudy (anonymous):

you are given four points in the plane A=(6,-6), B=(11,5), C=(16,-5) and D=(20,8). The graph of the function f(x) consist of three line segments AB, BC, CD find the integral [20,6]f(x)dx by using rule for integrals and interpreting the integral in terms of sums and/or differences of areas of elementary figures.

OpenStudy (anonymous):

so you just have to find the area under each individual lines and sum it up.

OpenStudy (anonymous):

use f(x) = mx+b substitute points A and B to find the equation of line AB, substitute points B and C to find the equation of line BC and substitute points C and D to find equation of line CD

OpenStudy (anonymous):

I actual missed this lecture cause i was ill so i read the book and still dont understand how exactly to do this i know that i find slope, then that some how will give me base and height to find out the area

OpenStudy (anonymous):

each point is a set of [ f(x) , x ]

OpenStudy (anonymous):

well. I gtg now lemme know if this works out otherwise I will post a full solution later.

OpenStudy (anonymous):

really confused i go AB=(11/5)x+5 BC=(-5/3)x-(40/3) and CD=(13/4)x+47

OpenStudy (anonymous):

I'm afraid I'm not going to be much help here. I need to brush up on my surface integrals.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

That was last year ;)

OpenStudy (anonymous):

Now it's all about vector spaces.

OpenStudy (anonymous):

oh ok lol

OpenStudy (anonymous):

\[\int\limits_{2}^{8}4xdx\] can you help with this one